Suppose an index of small company stocks started in 1956 at 10, and the index level was 1760.31 in 2016. What is the capital gains yield of the small company stocks for the period?
Answer #1

The capital gains yield = (Index level at 2016) - (Index level at 1956) / (Index level at 1956)

The capital gains yield = (1760.31 - 10 ) / 10 = 1750.31 / 10 = 175.031 or 17503.10%

The average annual capital gain yield = 17503.10% / 70 years = 250%